the capacitance of a spherical condenser is 1 micro F.if the spacing between two spheres is 1mm, the radious of the outer sphere is

Grade:12

2 Answers

arpit juneja
6 years ago
C=4pi*epsilon      1
      (1/ri)-(1/ro)
ATQ-   ro-ri=0.001m ,C=0.000001F      2
putting 2 in 1
therefore ro^2 -ro(0.001) -(9*1000000000*0.000000001) = 0
therefore ro = 3 metres(approx)
 
 
 
ankit singh
3 years ago

Capacitance of spherical conductor =4πϵ0a
where a is radius of conductor.
Therefore, C=9×1091×1=91×109
=0.11×109F=1.1×1010F